OpenStudy (anonymous):

classify all critical points g(x, y) = 1 − x^(2) − x − 2y^(2) + y

OpenStudy (anonymous):

Find where the gradient of g

OpenStudy (anonymous):

The gradient is {-1 - 2 x, 1 - 4 y} set it up ={0,0} and solve for x and y

OpenStudy (anonymous):

evaluating the first and second order partial derivatives would be a good start

OpenStudy (anonymous):

You find \[ \left\{\left\{x\to -\frac{1}{2},y\to \frac{1}{4}\right\}\right\} \]

OpenStudy (anonymous):

The hessian determinant is always 8. Now study the sign of the second derivative with respect to x

OpenStudy (anonymous):

The second derivative is always equal to -2, so (-1/2, 1/4) is a local max

OpenStudy (anonymous):

Here is a plot

OpenStudy (anonymous):

Are you still with me?

OpenStudy (anonymous):

yeah wait why when its X alone does it just stay as 1?

OpenStudy (anonymous):

meant to delete why

OpenStudy (anonymous):

I did not understand your question

OpenStudy (anonymous):

where did you get -1 from in -1-2x

OpenStudy (anonymous):

The partial derivative of g with respect to x

OpenStudy (anonymous):

so X equals 1? it doesn't just get cancelled out?

OpenStudy (anonymous):

what is the derivative of − x^(2) − x with respect to x It is -2 x -1

OpenStudy (anonymous):

okay that makes sense, i was a little confused if x = -1

OpenStudy (anonymous):

You might review your differentiation skills. You can do that on my site http://www.saab.org/calculus.cgi

OpenStudy (anonymous):

when solving for the partial derivative

OpenStudy (anonymous):

Do you understand it now?